OpenStudy (anonymous):

-18+4+x=9-19 what's the solution to the equation?

OpenStudy (anonymous):

is it 4

OpenStudy (anonymous):

x = 24 . do you need to know why?

OpenStudy (anonymous):

i do

OpenStudy (anonymous):

please? im so lost

OpenStudy (anonymous):

sorry hold up and quick sec

OpenStudy (whpalmer4):

Solution is easy to check: plug it into formula and see if you get a true statement.

OpenStudy (anonymous):

-18+4+x=9-19 -14+x=-10 +14 to the otherside 14-10 4

OpenStudy (whpalmer4):

@sunnybee1652 might want to check your work...

OpenStudy (anonymous):

yeah. mine is wrong. sorry about that.

OpenStudy (anonymous):

still not following geez louise *falls on floor*

OpenStudy (anonymous):

so its 4?

OpenStudy (anonymous):

answer is 4

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

was it right

OpenStudy (anonymous):

didn't check it yet im over whelmed with math I just don't understand it no matter how much I read and study this book

OpenStudy (anonymous):

so u say -18+4 first that's how you got -14?